In an experiment involving a treatment applied to 5 test subjects, researchers plan to use a simple random sample of 5 subjects selected from a pool of 7 available subjects. How many different random samples are possible???
Answer by DrBeeee(684)

7*6*5*4*3 = 2520
This results from;
any of the 7 for the first choice
only 6 for second, etc.
for 5 selections.
Note that there are 5 numbers multiplied together, one for each selection.
This has a name, "the of Permutations of 7 items selected 5 at a time, when order counts."
However, re-reading the experiment, the order of the 5 selected inividuals probably doesn't matter. This reduces the number by a factor of 5! or 5*4*3*2*1 = 120
The answer I think you want is 2520/120 = 21
This is called the number of "Combinations of 7 items selected 5 at a time."
Answer: There are 21 random samples of 5 victims.